What volume of 1.00 M NaOH is required to neutralize the following solution?

15.9 mL of 0.138 M phosphoric acid, H3PO4
1
Expert's answer
2019-10-07T06:42:38-0400

15.9 x 0.138/1000= 0.0021942 moles H3PO4.

3NaOH + H3PO4 = Na3PO4 + 3H2O

3:1=x:0.0021942

x= 0.0065826 moles of NaOH.

V= n/Cm = 0.0065826/1= 0.0065826 L= 6.5826 ml.